OpenStudy (anonymous):

Hi guys, can you explain to me the steps to solve this (2)/(5x+1)=(1)/(2x-1)?

geerky42 (geerky42):

Cross-multiply.

OpenStudy (anonymous):

2/( 5x+1)=1/(2x-1) 2(2x-1)=1( 5x+1) 4x-2= 5x+1 x= -3